OpenStudy (anonymous):

Add. Simplify, if possible. (4a)/(2a-4)+(8a)/(4a-8)

OpenStudy (anonymous):

i got (4a)/(a-2) is this right?

jhonyy9 (jhonyy9):

4a/(2(a-2)) +8a/(4(a-2)) = (8a +8a)/(4(a-2)) = 16a/(4(a-2)) = 4a/(a-2)

jhonyy9 (jhonyy9):

yes your answer is right